Lightning strikes can involve current upto 25000A and last upto 40millisec. If a person is struck by lightning with these properties, the current will pass through his body. We shall assume that his mass is 75 kg and that he is all wet with resistence 1.0 kohm. By how many degrees cecius would this lightning increase the temp og 75 kg of water. Given that internal temp of the body is 37 degreeC, would the temp increase that much ? Why not ? What would happen first ?

Grade:10

1 Answers

Sher Mohammad IIT Delhi
9 years ago
Mdt= i^2?*r*t
M(t-37)=25000*1000*40/1000
Depending upon the clothes etc , the resistance will be different.